### v1

# Quantized topological transport mediated by the long-range couplings

Abstract
Certain topological systems with time-varying Hamiltonian enable quantized and disorder-robust transport of excitations. Here, we introduce the modification of the celebrated Thouless pump when the on-site energies remain fixed, while the nearest and next-nearest neighbor couplings vary in time. We demonstrate quantized transport of excitations and propose an experimental implementation using an array of evanescently coupled optical waveguides.
Topological physics uncovers promising approaches to control the localization and propagation of light by tailoring the bandstructure of the material and harnessing localized or propagating topological modes [1, 2, 3].
Even richer physics arises when the Hamiltonian of the system varies in time. If such variation is periodic, this gives rise to Floquet physics and non-equilibrium phases with tailored properties [4, 5]. Of special interest is a periodic variation of the Hamiltonian resulting in the transport of excitations analogously to the water flow driven by the Archimedean screw. Leveraging the topological nature of the system such transport can be made disorder-resilient in the sense that the charge transferred during a single pumping cycle is robustly quantized [6, 7].
Historically the first example of such topological transport was the so-called Thouless pump [8] which utilized the Rice-Mele model [6, 7] with time-varying on-site energies and couplings between the sites. The latter system, in turn, is a generalization of the celebrated Su-Schrieffer-Heeger model [9], a paradigmatic example of a one-dimensional topological system.
The Thouless pump has been realized experimentally for ultracold atoms in a dynamically controlled optical lattice [10, 11, 12] as well as for photonic systems [13, 14, 15, 16]. Recent generalizations include systems with strong nonlinearity [17, 18, 19] and pumping of multiband systems [20, 21, 22] which could feature non-Abelian physics [20, 21]. Importantly, the protocol of Thouless pumping in simple two-band systems requires a synchronized variation of both couplings and on-site energies, which is quite challenging to implement in the arrays of evanescently coupled optical waveguides requiring a simultaneous modulation of the refractive index contrast and distance between the waveguides [16].
In this Letter we address that challenge and design an alternative protocol of topological pumping which only requires time-varying couplings. To compensate for the lack of time-varying on-site energies, we introduce next-nearest neighbor couplings. Below, we analyze this pumping scheme and demonstrate that it can be readily implemented using the arrays of laser-written evanescently coupled optical waveguides without the need to modulate their refractive index along the direction of propagation. In addition, we outline a prospective realization of this physics for another material platform – polariton condensates.
Specifically, we consider a one-dimensional lattice with the nearest-neighbor couplings $J_{1}$ and $J_{2}$ resembling the celebrated Su-Schrieffer-Heeger model. In addition, the structure is supplemented by the next-nearest neighbor couplings $t_{1}$ and $t_{2}$ connecting the sites of the same sublattice [Fig. 1(a)]. Hence, Bloch Hamiltonian of the periodic system is presented in the form
$H(k)=\begin{pmatrix}2t_{2}\cos k&J_{1}+J_{2}e^{-ik}\\ J_{1}+J_{2}e^{ik}&2t_{1}\cos k\end{pmatrix}\>$. (1)
To realize pumping, the Hamiltonian varies in time $\tau=t/T$, $T$ being the period of the pumping cycle, as follows:
$J_{1}(\tau)$ $=J_{0}-\frac{A}{2}\cos(2\pi\tau)\>,$ (2)
$J_{2}(\tau)$ $=J_{0}+\frac{A}{2}\cos(2\pi\tau)\>,$ (3)
$t_{1}(\tau)$ $=B\sin(2\pi\tau)\>,$ (4)
$t_{2}(\tau)$ $=-B\sin(2\pi\tau)\>,$ (5)
where $J_{0}$, $A$, and $B$ are constant factors defining the modulation amplitude. Figure 1(b) illustrates the dependence of the couplings $J_{1,2}$, $t_{1,2}$ on time. While all couplings vary harmonically, there is a $\pi/2$ phase shift between the nearest and next-nearest neighbor couplings. In addition, nearest-neighbor couplings $J_{1,2}$ remain positive throughout the entire pumping cycle, while the next-nearest neighbor couplings $t_{1,2}$ switch their sign.
Inspecting Bloch Hamiltonian Eq. (1), we observe that it is quite similar to that in the standard Thouless pump protocol. A formal difference appears in $\cos k$ term at the diagonal. However, this leads to a quite different physical realization which we investigate below.
Figure 1: (a) Schematic of the studied 1D lattice. Synchronized variation of nearest and next-nearest-neighbor couplings enables quantized topological transport. Dashed rectangle shows the unit cell including the sites of A and B sublattices. (b) Evolution of the couplings during one pumping cycle. Parameters: $J_{0}=1$ , $A=1$, $B=0.35$.
While the time-varying Hamiltonian has no stationary states, it is instructive to diagonalize it at the arbitrary moment of time $\tau$ and evaluate its instantaneous spectrum given by the expression
[…]
where the couplings $J_{1,2}$ and $t_{1,2}$ depend on time according to Eqs. (2)-(5).
The spectrum for a finite 14-unit-cell lattice in Fig. 2(a) provides an intuition on how the system behaves in the adiabatic limit, i.e. when the driving frequency is much smaller than the characteristic eigenfrequencies of the system. The results suggest that the gap in the spectrum remains open throughout the entire pumping cycle, and there are only two edge states which cross the bandgap and are responsible for the quantized transport. In Figure 2(a) the energies of left- and right-localized edge states are shown by red and blue colors, respectively.
In addition to the instantaneous spectrum we also examine the instantaneous eigenvectors. While Bloch functions do not provide immediate insights into the properties of the pumping scheme, their linear combinations known as Wannier functions can be readily visualized; they feature good localization, form an orthogonal set and are related to each other via translation [7].
To construct the Wannier functions, we first introduce the position operator in the form suitable for the periodic systems [23, 24, 25]
$\hat{X}=e^{\frac{2\pi i}{N}\hat{x}}$, (7)
where $\hat{x}$ yields the number of the unit cell $1,2,\dots N$. The operator $\hat{X}$ defined in this way respects the periodic boundary conditions, while the expectation values of the particle position are computed from its eigenvalues $X_{m}$ as $iN/(2\pi)\,\log(X_{m})$.
Figure 2: Instantaneous spectrum and Wannier centers of the driven system. (a) Evolution of the instantaneous spectrum calculated for a finite system consisting of $N=14$ unit cells. Left- and right-localized edge states are highlighted by red and blue, respectively. (b) Evolution of the instantaneous Wannier centers for the same system during a single driving cycle.
Next we project this position operator onto the lowest band using the projector
$\hat{P}=\sum_{n=1}^{N/2}|\Psi_{n}\rangle\langle\Psi_{n}|\>,$ (8)
[…]
$\hat{X}_{p}=\hat{P}\hat{X}\hat{P}\>$. (9)
The eigenvectors of the projected position operator $\hat{X}_{p}$ provide the Wannier functions of the lowest band [7]. At the same time, the eigenvalues of $\hat{X}_{p}$, $\lambda_{m}$, define so-called Wannier centers via
[…]
We plot the instantaneous Wannier centers for our system in Fig. 2(b) and observe that all Wannier centers shift by one unit cell during the pumping cycle. Since an arbitrary input state from the lowest band can be decomposed into the superposition of Wannier functions, this suggests that the designed protocol transports any such state by one unit cell during a single pumping cycle.
Figure 3: (a) Diffraction of the point-like wavepacket projected onto the lower Bloch band for the 14 unit cell lattice. (b) Displacement of the wavepacket center of mass by $1.005$ unit cell over one cycle $T=7$ (black) and corresponding Wannier center trajectory (red dashed).
So far we discussed the adiabatic limit of our model. To assess how the transport protocol works at a finite driving frequency, we solve the temporal evolution of the system consisting of $N=14$ unit cells. We choose the driving period $T=7$, which substantially exceeds the average inverse bandgap size $1/\Delta\simeq 0.6$, justifying the adiabatic approximation. We assume that the initial state is localized in the middle of the array and belongs entirely to the lowest band, i.e.
[…]
Solving the Schrödinger equation, we recover the probability distribution in Fig. 3(a) which shows that initially localized wavepacket starts to spread over the lattice. While the distribution $|\psi_{n}(t)|^{2}$ itself does not exhibit clear signatures of transport, it is instructive to track the motion of the wavepacket center of mass, i.e. $\left<x(t)\right>=\left<\psi\left|\hat{x}\right|\psi\right$> [Fig. 3(b)]. We observe that the center-of-mass motion resembles the temporal dependence of the Wannier centers and exhibits a shift equal to 1.005 unit cell during a single pumping cycle. Nearly quantized change of $\left<x\right$> is due to the topological nature of the pump, while slight violation of the quantization appears due to the non-adiabatic evolution of the system. Note that in order to have strictly quantized pumping at a finite driving frequency the temporal profile of the driving has to be fine-tuned [26] using, for instance, counter-adiabatic driving method [27, 28].
Figure 4: (a) Chosen geometry of the lattice and its spatial modulation. Insets show electric field profiles (amplitude) of symmetric and antisymmetric modes in vertical and detuned horizontal waveguides, respectively. (b) Modulation of the couplings $J_{1,2}$, $t_{1,2}$ and detunings $u_{1,2}$ extracted from numerically simulated splitting of the dimer eigenmodes. (c) Instantaneous spectrum of the propagation constants computed for a finite lattice with $14$ unit cells. Left- and right-localized edge states are highlighted by red and blue, respectively. (d) Tight-binding simulation of the real-space discrete diffraction pattern for the initial point-like excitation projected onto the lowest Bloch band of a $14$ unit cell lattice and modulation period $L=50$ cm. (e) Time dependence of the center of mass of the intensity distribution (black solid line), displaced by $1.0068$ unit cells over one modulation period. The trajectory of the Wannier center is shown by the dashed red curve for comparison.
As a specific platform to implement our protocol of topological pumping at optical wavelengths, we propose an optical waveguide lattice fabricated via the femtosecond laser writing technique [29, 30, 31]. We exploit a formal analogy between the tight-binding Schrödinger equation and the coupled-mode description of the waveguide lattice: the time variable $t=\tau T$ in the former corresponds to the propagation distance $z=\tau L$ in the waveguide system, where $L$ is the spatial period of the modulation along the waveguides.
We propose the zigzag-like geometry shown in Fig. 4(a) consisting of the main vertical waveguides shown in white and detuned connector waveguides depicted in red. The positioning and orientation of connector waveguides is chosen in such a way that they efficiently couple to the sublattice of main waveguides close to them, but remain practically decoupled from another sublattice of main waveguides. Depending on parameters, the waveguides can support the modes with the different symmetry of the near field profile. While horizontally oriented connector waveguides could utilize symmetric ($s$) mode, antisymmetric ($p_{x}$) modes odd under $\sigma_{v}$ reflection generally enable larger bandgap (see Supplementary Materials). Accordingly, we choose the refractive index contrast of connector waveguides such that the $s$-like modes of the vertical waveguides have the propagation constant closer to the $p$-like modes of the connector sites, thus realizing an instance of multi-orbital physics [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Alternatively, for the connector $p$-mode sites one could use fine-tuned photonic molecules based on pairs of closely placed vertically oriented waveguides [36].
Since the connector waveguides are assumed to be detuned from the main ones, they realize non-resonant coupling and can be excluded from the description using the degenerate perturbation theory [48]. By doing so, we recover an effective Hamiltonian possessing next-nearest-neighbor couplings $t_{1,2}=\kappa+\gamma^{2}/(\kappa+\Delta)$ between the main waveguides belonging to the same sublattice. In addition, main waveguides also acquire additional detunings $u_{1,2}=-\gamma^{2}/(\kappa+\Delta)$, assuming $|\Delta|\gg\gamma$. Here, $\Delta=k_{z}^{p}-k_{z}^{s}$ is the detuning in the propagation constants of $p$-modes in horizontal connector waveguides ($k_{z}^{p}$) from vertical $s$-modes in main lattice waveguides ($k_{z}^{s}$), $\gamma$ is the absolute value of the coupling between main and detuned waveguides, and $\kappa$ is the direct coupling between the waveguides belonging to the same sublattice. Importantly, the antisymmetric shape of $p$-like modes ensures that the couplings $t_{1,2}$ and detunings $u_{1,2}$ are modulated out of phase. As a result, the two mechanisms of bandgap opening add up, lead to the larger bandgap and allow to reduce the modulation period.
[…]
Such bending of the waveguides gives rise to the $z$-dependent couplings depicted in Fig. 4(b). Here, the effective couplings $J_{1,2}$, $t_{1,2}$, $\kappa$ and detunings $u_{1,2}$ are recovered by computing the eigenmodes of a pair of waveguides and tracking the splitting between the modes, which quantifies the strength of the coupling. The obtained values of the couplings justify the use of the degenerate perturbation theory for the chosen detuning $\Delta$.
Next we simulate numerically the spectrum for a lattice consisting of $14$ unit cells [Fig. 4(c)] and observe a good agreement with the tight-binding result using numerically calculated couplings (see Supplementary Materials for details). The spectrum shows left-and right-localized topological edge states traversing the complete bandgap in the opposite directions, indicating the presence of the quantized pump. Moreover, the spectrum in this case features a particularly large bandgap, facilitating quantized transport for smaller lattices and shorter modulation periods with the greater robustness to the disorder.
Finally, the tight-binding simulation of topological pumping in Fig. 4(d) shows the asymmetric spreading of the initial point-like excitation projected onto the lower Bloch band for the array consisting of $14$ unit cells with the modulation period $L=50$ cm. The center-of-mass trajectory of the intensity distribution in Fig. 4(e) (black curve) shows quantized transport and follows quite closely the trajectory of the respective Wannier center (red dashed curve) with the total shift of $1.0068$ unit cells per modulation period. For smaller modulation periods, quantized nature of the transport is violated resulting, for instance, in the shift of $0.84$ unit cell for the modulation period $L=20$ cm. Note that the lattice periods required for quantized adiabatic transport may be further decreased for lattices operating close to the flat-band or all-bands-flat conditions [34], and also in the nonlinear pumping regime when the diffraction of the wave packet is suppressed by the nonlinearity [17].
In summary, we have proposed a topological pump allowing to transfer the quantum state by varying nearest-neighbor and next-nearest-neighbor couplings. This protocol provides an alternative to the celebrated Thouless pump and does not involve the change of on-site energies, which is especially suitable for the arrays of evanescently coupled optical waveguides. The proposed scheme is not limited to optical waveguides and can be readily generalized to the other physical platforms, e.g. to the polariton condensates trapped in a slowly varying periodic potential [49, 50] with a conceptual scheme elaborated in the Supplementary Materials.
See Supplementary Materials for the details of numerical simulations of the system with symmetric and anti-symmetric connector modes in an optical waveguide lattice, including lattice geometry and chosen modulation scheme, application of the degenerate perturbation theory for connector sites, evaluation of the effective couplings and detuning of connector waveguides, comparison of the tight-binding and numerical finite-lattice spectra and tight-binding simulations of quantized pumping as well the proposed implementation for the lattice of coupled polariton condensates.
We acknowledge Andrei Stepanenko and Anton Nalitov for valuable discussions. Theoretical models were supported by the Russian Science Foundation (Grant No. 24-72-10069). Numerical simulations were supported by Priority 2030 Federal Academic Leadership Program.
Author declarations
Conflict of interest
The authors have no conflicts to disclose.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.